HOW CAN I SALVE THE PROBLEM -8X-5=3(6X-2)
Answer by checkley77(12844)

-8X-5=3(6X-2)
-8X-5=18X-6
-8X-18X=-6+5
-26X=-1
X=-1/-26
X=1/26 ANS.